OpenStudy (jacksonjrb):

Need help with a radical equation. Open question for equation. Will give medals.

OpenStudy (jacksonjrb):

\[\sqrt{2x+9}-\sqrt{x+1}=\sqrt{x+4}\]

OpenStudy (anonymous):

square both sides then you will have to isolate the radical and do it again it is a pain

OpenStudy (jacksonjrb):

whats would it look like after you squared it. thats all i need

OpenStudy (anonymous):

first off another easy inspection tells you the answer is zero when you square you get \[2x+9-2\sqrt{2x+9}\sqrt{x+1}-(x+1)=x+4\]

OpenStudy (anonymous):

combine like terms put everything without the radical on the other sides then square again

OpenStudy (jacksonjrb):

ok thanks.

OpenStudy (anonymous):

yw don't forget to square the \(-2\)

OpenStudy (jacksonjrb):

I'm sorry. @satellite73 Could you please do the second step? Then I think I'll be good.

OpenStudy (anonymous):

\[2x+9-2\sqrt{2x+9}\sqrt{x+1}-(x+1)=x+4\] distribute the \(-\)sign \[2x+9-2\sqrt{2x+9}\sqrt{x+1}-x-1=x+4\] combine like terms \[x+8-2\sqrt{2x+9}=x+4\]

OpenStudy (anonymous):

it is late and i messed up the last one \[x+8-2\sqrt{2x+9}\sqrt{x+1}=x+4\]

OpenStudy (jacksonjrb):

@satellite73 Sorry again. :P One more step then I'm almost certain I'll be good.